# Wormhole geometries in Einstein-aether theory

### v2

@@ 1 Introduction
The ringdown properties of black holes in Einstein–æther theory have been investigated, revealing that quasinormal modes are modified: damping rates generally decrease and ringing periods increase relative to GR, offering possible, but subtle, opportunities for gravitational-wave probes [50, 51]. Studies of quantum tunneling at both Killing and universal horizons in Einstein-Maxwell-æther black holes indicate nontrivial effects on Hawking radiation, especially under higher-order curvature corrections and modified dispersion relations [52].
@@ 2 Einstein-Aether theory and Field equations
For a static and spherically symmetric spacetime
@@
the field equations can be solved for certain choices of the coupling constants $c_{i}$’s [56]. In particular, the following classes of solutions have been identified for black holes. I: $c_{2}\neq 0$, $c_{13}\neq 0$, and $c_{14}=0$, II: $c_{2}\neq 0$, $c_{13}=0$, and $c_{14}=0$, III: $c_{2}=0$, $c_{13}\neq 0$, and $c_{14}=0$, IV: $c_{2}=0$, $c_{13}=0$, and $c_{14}\neq 0$, $\cdots$. In the next section, we focus on traversable wormhole solutions in EA theory and analyze their energy conditions.
3 Treversable wormhole solutions
Let us start with the line element of a traversable wormhole as in [5]
@@ 3 Treversable wormhole solutions
To preserve the metric signature throughout the wormhole spacetime $r>r_{0}$, it is also necessary that
@@
4 Class I: $c_{13}=0$, $c_{14}=0$, and $c_{2}\neq 0$
By choosing the coupling constants of the EA theory in this manner, the equations (3)-(3) become solvable, yielding [56]
@@ 4.1 Wormholes with power law shape function
At first glance, the wormhole solutions in this class depend on four parameters: $d$, $n$, $c_{2}$, and $r_{0}$. By setting $r_{0}=1$, we fix the unit of the radial coordinate $r$ to the size of the wormhole throat, leaving three parameters $d$, $n$, and $c_{2}$ in the solutions. In the following, we study the effect of these parameters on the energy conditions of the wormhole solutions.
In Fig. 1, the blue regions indicate areas in the parameter space of the wormhole solutions of the EA theory where ${\rm EC1}>0$, ${\rm EC2}>0$ and ${\rm EC3}>0$ simultaneously. In other words, for appropriate values of the parameters, the EA wormholes satisfy the WEC and, accordingly, the NEC. In particular, near the wormhole throat ($r_{0}=1$), the WEC and NEC are satisfied if $c_{2}>0$. Considering the constraints (2.7) on the coupling constants $c_{i}$ and noticing that $c_{13}=0$, one obtains $c_{2}>-\frac{2}{3}$, which is respected in our figures. However, our results indicate that the requirement of satisfying the energy conditions near the wormhole throat imposes a stronger constraint: $c_{2}>0$.
@@
Figure 1: The blue regions in these figures indicate where ${\rm EC1}>0$, ${\rm EC2}>0$, and ${\rm EC3}>0$ are satisfied simultaneously, meaning that the weak and null energy conditions are respected for the wormhole solutions in the EA theory. In these figures, we set $r_{0}=1$ and $d=5$. Fig. (a) shows that, by varying the coupling $c_{2}$, the WEC is satisfied near the wormhole throat if $c_{2}>0$ with $n=1$. In Fig. (b) the variation of the exponent $n$ versus $c_{2}$ is depicted for the radial distance $r=1.5$, and it is evident that the WEC is also satisfied for $c_{2}>0$.
In Fig. 2, by setting $c_{2}=\frac{1}{2}$, we show that the NEC (panel (a)) and WEC (panel (b)) are satisfied over the radial coordinate $r$ if the exponent $n$ in (4.3) lies within the interval indicated by the blue regions. A narrow gap appears around $r=2.5$, where neither the NEC nor WEC is satisfied. From Fig. 1, it is evident that in this narrow gap, the energy conditions are respected if $c_{2}$ takes negative values. Ignoring this small interval, one can conclude that the NEC and WEC are satisfied for the wormhole solutions of the EA theory when $c_{2}$ is positive. We also verified that variations in the parameter $d$ do not significantly affect these results.
(a)(b)
Figure 2: Satisfaction of the NEC and WEC for arbitrary values of the radial coordinate $r$ in the wormhole solutions of the EA theory is illustrated in these figures, plotted for $c_{2}=\frac{1}{2},d=5$, and $r_{0}=1$. In (a), the blue regions indicate where ${\rm EC2}>0$ and ${\rm EC3}>0$, showing that the NEC is satisfied for appropriate values of $n$ over nearly all radial distances outside the wormhole throat. A small gap appears around $r=2.5$ where the NEC is violated; however, within this gap, the NEC is satisfied for negative values of $c_{2}$ in the range $\frac{2}{3}<c_{2}<0$. A similar analysis is presented in (b) for the WEC, which is satisfied over a more limited region compared to the NEC.
@@ 4.2 Wormholes with logarithmic shape function
Noting the conditions (3.2)-(3.4), it is straightforward to show that the wormhole throat in this case must satisfy the bound $r_{0}>1$. By substituting the shape function (4.5) into (4), one can then obtain the corresponding expressions for the energy conditions as
@@
To investigate the WEC and NEC, we examine whether there exist regions in the parameter space of these wormhole solutions where the above energy conditions are positive. Fig. 3 illustrates this parameter space. The blue regions in the figure indicate where EC1>>0, EC2>>0, and EC3>>0 simultaneously. It is evident from Fig. 3 that the NEC and WEC are violated for $c_{2}<0$. However, for suitable positive values within the bound $c_{2}>0$, both the NEC and WEC are satisfied at the wormhole throat and at large distances from it (except for a narrow interval). In other words, the satisfaction of the NEC and WEC for wormholes with the logarithmic shape function imposes a constraint on the coupling constant of the EA theory: $c_{2}>0$.
@@
Figure 3: Satisfaction of the energy conditions for arbitrary values of the radial coordinate $r$ in the case of wormhole solutions with a logarithmic shape function in the EA theory. The blue regions in (a) indicate where ${\rm EC2}>0$ and ${\rm EC3}>0$ simultaneously, implying that the NEC is satisfied. In (b), the blue regions correspond to the domain where all EC1, EC2, EC3 are positive, meaning that the WEC is respected. Our analysis shows that, for appropriate positive values of $c_{2}>0$, both the NEC and WEC are satisfied for nearly all radial distances outside the wormhole throat. These figures are plotted for $r_{0}=2$; however, we have verified that varying $r_{0}$ (within the bound $r_{0}>1$) does not significantly affect this overall behavior.
@@ 4.3 Wormholes with hyperbolic shape function
We performed a similar analysis and identified regions in the parameter space of these wormhole solutions where the energy conditions are satisfied. The results are presented in Fig. (4). As seen in this figure, when the EA coupling constant satisfies the bound $c_{2}>0$, the NEC is respected both near the wormhole throat and at large radial distances. In part (b) it is also evident that the WEC is satisfied around the wormhole throat for $c_{2}>0$.
@@
Figure 4: The wormhole solutions with the hyperbolic shape function satisfy the NEC (panel a) and WEC (panel b) in the blue regions. By choosing a suitable value of the coupling constant of the EA theory such that $c_{2}>0$, the NEC is satisfied both at the wormhole throat and at large radial distances. The WEC is also respected around the wormhole throat for appropriate values within the bound $c_{2}>0$. In these figures we set $r_{0}=2$; however varying $r_{0}$ within the bound $r_{0}>0$ does not alter this overall behavior.
@@ 5 Class II: $c_{2}=0$, and $c_{14}=0$
At this stage, we substitute three types of shape functions, power-law, logarithmic, and hyperbolic, into the above equations and investigate the corresponding energy conditions for the wormhole solutions of the EA theory in this case.
@@ 5.1 The power law shape function
Note that by setting $r_{0}=1$, only two parameters $e$ and $n$ and the combination $c_{3}-c_{4}$ remain for the solutions in this class. We are now in a position to investigate the energy conditions and identify regions in the parameter space where the NEC and WEC are satisfied. Due to the constraints (3.2)-(3.4) the parameter $n$ in the power-law shape function (4.3) should satisfy the bound n>1n>-1, which has been taken into account in our analysis.
Figure 5: The inequalities ${\rm EC1}>0$, ${\rm EC2}>0$, and ${\rm EC3}>0$ are satisfied simultaneously in the blue regions, indicating that both the WEC and NEC are respected for the wormhole solutions with the power-law shape function in the EA theory. Here, we set $r_{0}=1$, $n=1$, and $e=3$. It is evident that, except for a narrow interval around $r=2$ (twice the throat radius), the WEC and NEC are satisfied throughout the spacetime for the wormhole solutions of the EA theory, provided that the combination $c_{3}-c_{4}>0$ .
The blue region in Fig. 5 indicates where ${\rm EC1}>0$, ${\rm EC2}>0$ and ${\rm EC3}>0$ simultaneously, meaning that the WEC and NEC are satisfied. From Fig. 5 it is evident that if the coupling constants $c_{3}$ and $c_{4}$ satisfy the bound $c_{3}-c_{4}>0$, the WEC is respected around the wormhole throat (located at $r_{0}=1$). Moreover, except for a narrow interval around $r=2$, the WEC and NEC are satisfied for all remaining $r>r_{0}$, when $c_{3}-c_{4}$ is positive.
To explore the effect of the exponent $n$ on the WEC and NEC, we plot Fig. 6. It is clear from (a) that as $n$ decreases, the required value of $c_{3}-c_{4}$ increases. However, for any $n$ satisfying n>1n>-1, the WEC and NEC are preserved if $c_{3}-c_{4}>0.5$. As shown in part (b) of Fig. 6, by fixing $c_{3}-c_{4}=1$, the WEC and NEC are satisfied for almost all radial distances, except for a narrow range around $r=2$ for any exponent n>1n>-1.
We also performed a similar analysis for the satisfaction of $EC2>0$ and $EC>03$ (corresponding to the NEC only, where the WEC may be violated). The results show no significant differences. In other words, the wormhole solutions of the EA theory with $c_{2}=0$ and $c_{14}=0$ satisfy both the WEC and NEC if the combination $c_{3}-c_{4}$ satisfies the bound $c_{3}-c_{4}>\frac{1}{2}$. We also verified that varying the parameter $e$ does not significantly affect these results.
@@
Figure 6: The effect of $n$ on the satisfaction of the NEC and WEC in the wormhole solutions of the EA theory is illustrated in these figures. The plots are generated by setting $r_{0}=1,e=5$. Part (a) plotted for $r=3$ shows that for large values of $n$, the WEC and NEC are satisfied when $c_{3}-c_{4}>0$; as $n$ decreases, the required value of $c_{3}-c_{4}$ must increase. However, if $c_{3}-c_{4}>0.5$ the WEC and NEC are satisfied for all allowed values of $n$. In part (b) we sett $c_{3}-c_{4}=1$ and it is evident that the WEC and NEC are satisfied for almost all radial distances, independently of the value of $n$.
@@ 5.2 The logarithmic shape function
Here, the bound $r_{0}>1$ must be satisfied at the wormhole throat due to the conditions (3.2)-(3.4). Similar to the previous cases, we focus on solutions that satisfy the NEC and WEC. In Fig. 7 we depict the region in which all energy conditions in (5.2) are positive, indicating that the WEC is satisfied. It is evident that, for wormhole solutions with a logarithmic shape function, the WEC is satisfied at the throat and almost everywhere, provided appropriate values for the couplings within the bound $c_{3}-c_{4}>0$ are chosen.
@@
Figure 7: Wormhole solutions of the EA theory with a logarithmic shape function satisfy the WEC and NEC in the blue region. By choosing a suitable value for the coupling within the bound $c_{3}-c_{4}>0$, the WEC and NEC are satisfied at the wormhole throat (located at $r_{0}=2$) as well as at large distances. This figure is depicted for $e=5$; however, we verified that other values of $e$ and $r_{0}$ do not alter this overall behavior.

### v3

@@ 1 Introduction
The ringdown properties of black holes in Einstein–æther theory have been investigated, revealing that quasinormal modes are modified: damping rates generally decrease and ringing periods increase relative to GR, offering possible, but subtle, opportunities for gravitational-wave probes [50, 51]. Studies of quantum tunneling at both Killing and universal horizons in Einstein-Maxwell-æther black holes indicate nontrivial effects on Hawking radiation, especially under higher-order curvature corrections and modified dispersion relations [52]. The study of wormholes in Lorentz violating theories is also done in [53] where the authors obtained Exact traversable wormhole solution in bumblebee gravity.
@@ 2 Einstein-Aether theory and Field equations
In [57] the field equations of EA theory are solved for static and spherically symmetric spacetime
@@
In particular, the authors of [57], obtained some classes of black hole solutions of the EA theory, for certain choices of the coupling constants $c_{i}$’s as I: $c_{2}\neq 0$, $c_{13}\neq 0$, and $c_{14}=0$, II: $c_{2}\neq 0$, $c_{13}=0$, and $c_{14}=0$, III: $c_{2}=0$, $c_{13}\neq 0$, and $c_{14}=0$, IV: $c_{2}=0$, $c_{13}=0$, and $c_{14}\neq 0$, $\cdots$. In the next section, we focus on traversable wormhole solutions in EA theory and analyze their energy conditions.
3 EA traversable wormhole geometries
Teraversable wormhole solutions are introduced in [5]. In this article we consider the traversable wormholes with vanishing tidal force, in which the line element is in the form
@@ 3 EA traversable wormhole geometries
To preserve the metric signature throughout the wormhole spacetime, the radial coordinate satisfies the bound $r>r_{0}$. It is also necessary that
@@
4 Class I: $c_{13}=0$, $c_{14}=0$ and $c_{2}\neq 0$
By choosing the coupling constants of the EA theory as $c_{13}=0,c_{14}=0$ and $c_{2}\neq 0$, the equations (3)-(3) become solvable, yielding [57]  [^4] In [57] the equations of motion of the EA theory are solved for static spherically symmetric black hole metric (2.13) and the authors find metric functions $A(r)$ and $B(r)$ and the function $a(r)$ of the aether field. In this work, we are going to study the wormhole geometries (for which the metric is in spherically symmetric form) in the background of EA theory. So we encounter the same equations for the aether field as in [57] and we use their solutions.
@@ 4.1 Wormholes with power law shape function
The above terms possess singularity at $r=\sqrt{d}$ which are eliminated by setting $\sqrt{d}<r$. Remember that radial coordinate in the wormhole space-time satisfies $r_{0}<r$, so if the integration constant $d$ respects the bound $\sqrt{d}<r_{0}$, the energy conditions are well-behaved throughout the wormhole space-time. In what follows, considering the bounds $\sqrt{d}<r_{0}$ and n>1n>-1, we investigate the parameter space of the wormhole solutions of the EA theory, in order to find regions in which the EC1, EC2 and EC3 in (4.1) take positive values.
In Fig. 1, we investigate the energy conditions around the wormhole throat (located at $r_{0}=\frac{1}{2}$). In this figure, the blue regions in panel (a) and (c) indicate areas where ${\rm EC1}>0$, ${\rm EC2}>0$ and ${\rm EC3}>0$ simultaneously. In other words, for appropriate values of the parameters, the EA wormholes satisfy the WEC and, accordingly, the NEC. The blue region in panel (b) also shows the satisfaction of ${\rm EC2}>0$ and ${\rm EC3}>0$, means that the NEC is respected. In particular, near the wormhole throat ($r_{0}=\frac{1}{2}$), the WEC and NEC are satisfied if $c_{2}>0$. Considering the constraints (2.7) on the coupling constants $c_{i}$ and noticing that $c_{13}=0$, one obtains $c_{2}>-\frac{2}{3}$, which is respected in our figures. However, our results indicate that the requirement of satisfying the energy conditions near the wormhole throat imposes a stronger constraint: $c_{2}>0$.
(a)(b)(c)
Figure 1: The blue regions in panels (a) and (c) indicate where ${\rm EC1}>0$, ${\rm EC2}>0$, and ${\rm EC3}>0$ are satisfied simultaneously, meaning that the WEC is respected for the wormhole solutions in the EA theory. The blue regions in panel (b) shows ${\rm EC2}>0$, and ${\rm EC3}>0$, means that NEC is respected. In these figures, we set $r_{0}=\frac{1}{2}$ and $d=0.2$. In Figs (a) and (b) by setting $n=-\frac{1}{2}$, we show that, by varying the coupling $c_{2}$, the WEC and NEC are satisfied near the wormhole throat if $c_{2}>0$. In Fig. (c) the effect of variation of $n$ versus $c_{2}$ on the satisfaction of WEC at the wormhole throat $r=r_{0}=\frac{1}{2}$ is depicted. It is evident that the WEC is also satisfied for $c_{2}>0$.
In order to explore the effect of exponent $n$ on the energy conditions, we have plotted Fig. 2. In this figure by setting $c_{2}=\frac{1}{2}$, we show that the NEC (panel (a)) and WEC (panel (b)) are satisfied over the radial coordinate $r$ if the exponent $n$ in (4.3) lies within the interval indicated by the blue regions. It is worth to mention that if one plot the same figure by setting $c=-\frac{1}{2}$ (or other negative values), the result is null. This means that the by varying $n$, NEC and WEC could not be satisfied if the coupling constant $c_{2}$ takes negative values. One can conclude that the NEC and WEC are satisfied for the power law shape function wormhole solutions of the EA theory when $c_{2}$ is positive.
@@
Figure 2: Satisfaction of the NEC and WEC for arbitrary values of the radial coordinate $r$ in the wormhole solutions of the EA theory is illustrated in these figures, plotted for $c_{2}=\frac{1}{2},d=0.2$, and $r_{0}=\frac{1}{2}$. In (a), the blue regions indicate where ${\rm EC2}>0$ and ${\rm EC3}>0$, showing that the NEC is satisfied for appropriate values of $n$ over nearly all radial distances outside the wormhole throat. A similar analysis is presented in (b) for the WEC, which is satisfied over a more limited region compared to the NEC.
@@ 4.2 Wormholes with logarithmic shape function
Noticing the conditions (3.2)-(3.4), it is straightforward to show that the wormhole throat in this case must satisfy the bound $r_{0}>1$. By substituting the shape function (4.5) into (4), one can then obtain the corresponding expressions for the energy conditions as
@@
Similar to (4.1), the above terms are singular at $r=\sqrt{d}$. However if the bound $\sqrt{d}<r_{0}$ is respected, then the singularity at $r=\sqrt{d}$ is eliminated from the wormhole space-time and energy conditions are well-behaved everywhere. Now to investigate the WEC and NEC, we examine whether there exist regions in the parameter space of these wormhole solutions where the above energy conditions are positive.
Fig. 3 illustrates this parameter space. It is evident from this figure that the NEC and WEC are violated for $c_{2}<0$. The blue regions in panel (a) of this figure indicate where EC2>>0 and EC3>>0 simultaneously which means the NEC is satisfied for suitable positive values within the bound $c_{2}>0$, on the wormhole throat and also at wide range of the radial coordinate $r$. In panel (b) the blue color means the simultaneous satisfaction of EC1>>0, EC2>>0 and EC3>>0. This shows that the WEC is respected outside the wormhole throat, if one choose the right values of the coupling constant within the bound $c_{2}>0$. In other words, satisfaction of the NEC and WEC for wormholes with the logarithmic shape function imposes a constraint on the coupling constant of the EA theory: $c_{2}>0$.
@@
Figure 3: Satisfaction of the energy conditions for arbitrary values of the radial coordinate $r$ in the case of wormhole solutions with a logarithmic shape function in the EA theory. The blue regions in (a) indicate where the NEC is satisfied. In (b), the blue regions correspond to the domain where all EC1, EC2, EC3 are positive, meaning that the WEC is respected. Our analysis shows that, for appropriate positive values of $c_{2}>0$, both the NEC and WEC are satisfied for nearly all radial distances outside the wormhole throat. These figures are plotted for $r_{0}=3$ and $d=8$.
@@ 4.3 Wormholes with hyperbolic shape function
We performed a similar analysis and identified regions in the parameter space of these wormhole solutions where the energy conditions are satisfied (once again we respect the bound $\sqrt{d}<r_{0}$ in order to omit the singularity at the wormhole spacetime). Our results show that the WEC could not be satisfied on the wormhole throat and beyond that (see panel (b) of Fig 4). However the NEC is respected at the wormhole throat and also at the large radial distances, if the EA coupling constant satisfies the bound $c_{2}>0$ (panel (a) in Fig. 4).
@@
Figure 4: The wormhole solutions with the hyperbolic shape function satisfy the NEC (panel a) in the blue regions. By choosing a suitable value of the coupling constant of the EA theory such that $c_{2}>0$, the NEC is satisfied both at the wormhole throat and at large radial distances. The absence of blue color in panel (b) shows that the WEC could not be respected for these wormholes. In these figures we set $r_{0}=5$ and $d=23$; however varying $r_{0}$ within the bound $r_{0}>\sqrt{d}$ does not alter this overall behavior.
@@ 5 Class II: $c_{2}=0$, and $c_{14}=0$
Similar to the previous section we encounter a singularity at $r=\sqrt{e}$ which eliminates if $\sqrt{e}<r_{0}$ and during this section, we respect this bound to obtain well-behaved solutions. At this stage, we substitute three types of shape functions, power-law, logarithmic, and hyperbolic, into the above equations and investigate the corresponding energy conditions for the wormhole solutions of the EA theory in this case.
@@ 5.1 The power law shape function
We are now in a position to investigate the energy conditions and identify regions in the parameter space where the NEC and WEC are satisfied. Due to the constraints (3.2)-(3.4) the parameter $n$ in the power-law shape function (4.3) should satisfy the bound n>1n>-1, which has been taken into account in our analysis.
Figure 5: The inequalities ${\rm EC1}>0$, ${\rm EC2}>0$, and ${\rm EC3}>0$ are satisfied simultaneously in the blue regions, indicating that both the WEC and NEC are respected for the wormhole solutions with the power-law shape function in the EA theory. Here, we set $r_{0}=\frac{3}{2}$, $n=-\frac{3}{4}$, and $e=2$. It is evident that, except for a gap, the WEC and NEC are satisfied throughout the entire spacetime of wormhole solutions, if the combination $c_{3}-c_{4}$ set adequately in the bound $c_{3}-c_{4}>0$ .
The blue region in Fig. 5 indicates where ${\rm EC1}>0$, ${\rm EC2}>0$ and ${\rm EC3}>0$ simultaneously, meaning that both WEC and NEC are satisfied. From Fig. 5 it is evident that if the coupling constants $c_{3}$ and $c_{4}$ are set in a manner that the combination $c_{3}-c_{4}$ take positive value, the WEC is respected on the wormhole throat (located at $r_{0}=\frac{3}{2}$). Moreover, the WEC and NEC are satisfied through a wide range outside the throat, when $c_{3}-c_{4}$ is positive.
To explore the effect of the exponent $n$ on the WEC and NEC, we plot Fig. 6. It is clear from (a) that as $n$ decreases, the required value of $c_{3}-c_{4}$ increases. However, for any $n$ satisfying n>1n>-1, the WEC and NEC are preserved if $c_{3}-c_{4}>0.5$. As shown in part (b) of Fig. 6, by fixing $c_{3}-c_{4}=5$, both WEC and NEC are satisfied at the wormhole throat if the exponent $n$ takes small values. Outside the wormhole throat, the WEC and NEC are also satisfied almost at all radial distances for any exponent n>1n>-1.
We also performed a similar analysis for the satisfaction of ${\rm EC2}>0$ and ${\rm EC3}>0$ (corresponding to the NEC only, where the WEC may be violated). The results show no significant differences. In other words, the wormhole solutions of the EA theory with $c_{2}=0$ and $c_{14}=0$ satisfy both the WEC and NEC if the combination $c_{3}-c_{4}$ satisfies the bound $c_{3}-c_{4}>\frac{1}{2}$. We also verified that varying the parameter $e$ does not significantly affect these results.
@@
Figure 6: The effect of $n$ on the satisfaction of the NEC and WEC in the wormhole solutions of the EA theory is illustrated in these figures. The plots are generated by setting $r_{0}=\frac{3}{2},e=2$. Part (a) plotted for $r=4$ shows that for large values of $n$, the WEC and NEC are satisfied when $c_{3}-c_{4}>0$; as $n$ decreases, the required value of $c_{3}-c_{4}$ must increase. However, if $c_{3}-c_{4}>0.5$ the WEC and NEC are satisfied for all allowed values of $n$. In part (b) we fix $c_{3}-c_{4}=5$. It is evident that WEC and NEC are respected at the throat for small values of $n$. However, outside the throat, WEC and NEC are satisfied for almost all radial distances, independently of the value of $n$.
@@ 5.2 The logarithmic shape function
Here, the bound $r_{0}>1$ must be satisfied at the wormhole throat due to the conditions (3.2)-(3.4). Similar to the previous cases, we focus on solutions that satisfy the NEC and WEC. In Fig. 7 we depict the region in which all energy conditions in (5.2) are positive, indicating that the WEC is satisfied. It is evident that, for wormhole solutions with a logarithmic shape function, the WEC is satisfied at the neighborhood of wormhole throat and almost throughout the entire space-time, provided values for the couplings within the bound $c_{3}-c_{4}>0$ are chosen.
@@
Figure 7: Wormhole solutions of the EA theory with a logarithmic shape function satisfy the WEC and NEC in the blue region. By choosing the coupling combination within the bound $c_{3}-c_{4}>0$, the WEC and NEC are satisfied at the wormhole throat (located at $r_{0}=1.01$) as well as at large distances. This figure is depicted for $e=\frac{1}{2}$.